OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (xapproachesinfinity):

well just plug in 3,4,5 to find the rest of the terms you given the first term 1 the second term 2 use those to find the rest

OpenStudy (xapproachesinfinity):

example f(3)=f(3-1)+f(3-2)=f(2)+f(1)=2+1=3 and so forth...